# Can MLLMs Guide Me Home? A Benchmark Study on Fine-Grained Visual Reasoning from Transit Maps

### v2

@@ Abstract
Multimodal large language models (MLLMs) have recently achieved significant progress in visual tasks, including semantic scene understanding and text-image alignment, with reasoning variants enhancing performance on complex tasks involving mathematics and logic. However, their capacity for reasoning tasks involving fine-grained visual understanding remains insufficiently evaluated. To address this gap, we introduce ReasonMap, a benchmark designed to assess the fine-grained visual understanding and spatial reasoning abilities of MLLMs. ReasonMap encompasses high-resolution transit maps from $30$ cities across $13$ countries and includes $1,008$ question-answer pairs spanning two question types and three templates. Furthermore, we design a two-level evaluation pipeline that properly assesses answer correctness and quality. Comprehensive evaluations of $15$ popular MLLMs, including both base and reasoning variants, reveal a counterintuitive pattern: among open-source models, base models outperform reasoning ones, while the opposite trend is observed in closed-source models. Additionally, performance generally degrades when visual inputs are masked, indicating that while MLLMs can leverage prior knowledge to answer some questions, fine-grained visual reasoning tasks still require genuine visual perception for strong performance. Our benchmark study offers new insights into visual reasoning and contributes to investigating the gap between open-source and closed-source models.
@@ 1 Introduction
Figure 1: Overview of ReasonMap. ReasonMap is a benchmark dataset designed to evaluate fine-grained visual reasoning abilities of MLLMs, encompassing $1,008$ question–answer pairs constructed over high-resolution transit maps from $30$ cities, spanning two question types and three templates.
@@
Multimodal large language models (MLLMs) [1, 2, 3, 4, 5] have recently achieved notable advancements across a range of vision-language tasks, including visual grounding [6, 7], reasoning segmentation [8, 9, 10, 11], and text-image alignment [12, 13]. Building upon these developments, reasoning MLLMs [14, 15, 16, 17, 18, 19, 20] have further improved performance on complex visual reasoning tasks such as visual math problems [21, 22], visual question answering (VQA) [23, 24, 22], and spatial reasoning [23, 25, 26]. These capabilities are critical for a wide range of real-world applications, including embodied AI, autonomous agents, and decision-making systems such as autonomous driving [27, 28, 29]. As multimodal tasks grow in complexity and practical relevance, the need for rigorous benchmarks to assess fine-grained visual reasoning becomes increasingly essential.
To address the growing demand for robust evaluation of multimodal reasoning, several benchmarks have been proposed. Datasets such as MathVQA [22] and MMMU [24] incorporate multimodal questions but often permit models to succeed via shallow heuristics, without requiring genuine visual grounding. MathVerse [30] mitigates this limitation by introducing diverse problem variants that encourage reliance on visual input. VisuLogic [31] further enforces visual reasoning by explicitly eliminating language-only shortcuts. Other efforts, such as VisualPuzzles [32], VGRP-Bench [33], and R-Bench [34], target logical and structural reasoning, while CityBench [35] and DriveBench [36] focus on domain-specific applications like urban tasks and autonomous driving. VBench [37] emphasizes detailed visual understanding. MapBench [38] addresses spatial reasoning by introducing structured scene graphs for map navigation. Despite these advances, systematic evaluation of fine-grained visual reasoning remains limited, especially for structured and information-rich diagrams like high-resolution transit maps, leaving a critical gap in existing benchmarks.
@@
In this paper, we introduce ReasonMap (Figure 1), a benchmark designed to evaluate the fine-grained visual understanding and spatial reasoning capabilities of MLLMs using high-resolution transit maps. As structured and information-dense visual artifacts, transit maps inherently require precise spatial interpretation, making them well-suited for assessing detailed visual reasoning. ReasonMap comprises $1,008$ human-verified question-answer pairs spanning $30$ cities across $13$ countries. Each instance includes a map, two stops, two questions (short and long), multiple reference routes, and two difficulty labels (map and question difficulty). The questions cover two types and three prompting templates—one for short and two for long questions—capturing both coarse and fine-grained spatial reasoning. To ensure data quality, we perform manual route verification, promote question diversity, and balance difficulty distribution. For evaluation, we propose a two-level framework that independently measures answer correctness (via accuracy) and quality (via a proposed map score), reflecting both feasibility and efficiency in route planning.
@@
We conduct comprehensive experiments on $15$ widely-used MLLMs, encompassing base and reasoning models. Our results reveal a counterintuitive finding: among open-source models, base variants outperform their reasoning counterparts, whereas the opposite holds for closed-source models. Moreover, when only textual inputs are provided, models can still answer some questions based on inner knowledge, but in most cases, their performance noticeably drops. This highlights a critical limitation in the current model behavior. While some models can leverage prior knowledge and textual cues to solve certain tasks, the tasks (e.g., fine-grained visual reasoning tasks) requiring genuine visual understanding still necessitate effective integration of multimodal information for robust reasoning.
Our main contributions are summarized as follows:
  • We develop an extensible, semi-automated pipeline for dataset construction, facilitating scalable expansion to additional maps and cities. Using this pipeline, ReasonMap is constructed to evaluate fine-grained visual reasoning capabilities in MLLMs.
  • We propose a structured two-level evaluation framework that separately quantifies answer correctness and quality using accuracy and the proposed map score, respectively, enabling fine-grained answer assessment.
  • A comprehensive benchmarking study is conducted across $15$ MLLMs, providing insights into model performance, robustness, and the interplay between visual and textual cues, thereby informing future research on multimodal reasoning.
@@ 2 Related Work
Reasoning in LLMs & MLLMs. Recent advances in large language models (LLMs) have demonstrated significant improvements in reasoning capabilities through reinforcement fine-tuning paradigms [14, 15, 39, 40], which leverage GRPO [41] to unlock the reasoning potential of LLMs. This paradigm has also been extended to the multimodal domain, with increasing interest in applying reinforcement learning (RL) to visual reasoning [42, 43, 44, 45, 46]. Both open-source and closed-source communities have introduced advanced reasoning MLLMs built upon earlier systems [3, 2, 47, 48]. Notable open-source models include Kimi-VL [16], Skywork-R1V [17, 18], and Qwen-QVQ [20], while Doubao-1.5-Pro [19], OpenAI o3 [48], OpenAI 4o [49], and Gemini [50] represent leading closed-source efforts. Despite recent progress, systematic evaluation of fine-grained visual reasoning in MLLMs remains limited, as existing benchmarks primarily target coarse-grained tasks and fail to capture model performance on complex real-world visual content.
Table 1: Comparison between ReasonMap and existing multimodal reasoning datasets. For entries in the dataset size column with notation like (×$\times n$), each base problem has multiple versions to enforce visual grounding. Specifically, VGRP-Bench is constructed by sampling over 20 core puzzles.
Name Year Dataset Size Avg. Resolution Training Set Step Evaluation Multilingual (Count)
MMMU [24] 2024 $11.5$k $684\times 246$ ($2$)
MathVerse [30] 2024 $2,612$ (×$\times 6$) $577\times 487$
VisuLogic [31] 2025 $1,003$ $601\times 331$
VisualPuzzles [32] 2025 $1,168$ $767\times 464$
VGRP-Bench [33] 2025 $20$ (×$\times 5$) $790\times 790$
R-Bench [34] 2025 $665$ $629\times 348$ ($2$)
@@
ReasonMap 2025 $1,008$ (×$\times 2$) $5,839\times 5,449$ ($4$)
Multimodal Reasoning Datasets. As multimodal reasoning has rapidly progressed, various benchmarks have emerged to evaluate MLLMs across different reasoning dimensions (see summary in Table 1). Datasets such as VBench [37], VisualPuzzles [32], VisuLogic [31], and VGRP-Bench [33] primarily examine abstract visual reasoning through synthetic tasks involving logic, structure, and pattern recognition. In parallel, CityBench [35] and DriveBench [36] shift focus to real-world spatial reasoning, assessing model performance on complex urban or autonomous driving scenarios. For mathematical reasoning, benchmarks like MathVQA [22], MMMU [24], and MathVerse [30] integrate multimodal inputs, with MathVerse notably introducing varied problem formats to strengthen visual dependence. Additionally, MapBench [38] employs structured scene graphs combined with CoT prompting to support navigation tasks based on manually curated and verified questions. Its image resolution, while relatively low, reflects a common characteristic shared by current datasets. Unlike these works, we first introduce a benchmark for evaluating fine-grained visual reasoning capacities with high-resolution transit maps.
Map-based Spatial Reasoning. Among the many directions of multimodal reasoning, map-based spatial reasoning has emerged as a crucial area, with broad applications in navigation, urban planning, and autonomous systems [51, 52, 53, 54]. Recent efforts have focused on enabling models to interpret and reason over various types of map data. CityBench [35] provides a dataset for evaluating urban scene understanding, while MapLM [55] introduces a benchmark for map and traffic scene understanding. PlanAgent [56] and PReP [57] explore embodied planning in environments that require interpreting map information. MapEval [26] proposes a structured evaluation suite for map-related reasoning, and GeoNav [53] investigates geospatial navigation using LLMs. Most existing methods [26, 35, 56] depend on external tools (e.g., map services or APIs) to complete spatial tasks, which often bypasses the need for genuine visual reasoning. However, spatial reasoning based solely on visual understanding remains essential. Our work targets to fill this gap by evaluating such capabilities without tool assistance.
@@ 3 ReasonMap Construction
In this section, we first present the complete dataset building pipeline as shown in Figure 2, which consists of three main stages: (1) data collection and preprocessing; (2) question–answer pair construction; and (3) quality control. We then report comprehensive statistics of the dataset.
@@ 3.1.1 Data Collection and Preprocessing
We collect and manually select $30$ high-resolution transit maps covering $30$ cities across $13$ countries from publicly available online sources, in compliance with relevant licenses and regulations, ensuring diversity and a balanced range of map difficulty. We then leveraged MLLMs to extract the names of transit lines and their corresponding stops, followed by manual correction to ensure correctness. Special cases such as transfer stops and branch-starting stops were annotated in a standardized format appended to the respective stop entries. Finally, for subsequent usage, all route and stop information was saved in a unified JSON format, referred to as the Map Metadata.
@@ 3.1.2 Question-Answer Pair Construction
The construction of question–answer pairs involves three key steps: (1) Question Generation, where we formulate questions based on predefined templates; (2) Reference Route Collection, where we obtain corresponding reference routes using Gaode Map [^1] https://console.amap.com/dev/index and Google Map [^2] https://developers.google.com/maps/apis-by-platform; and (3) Label Annotation, where we assign difficulty levels for both the maps and the questions.
@@
Label Annotation. Two levels of difficulty labeling are included in this stage. For map difficulty, we manually assign each map to one of three difficulty levels (easy, medium, hard), ensuring a balanced split across $30$ maps, with $10$ maps per level. For question difficulty, we assign difficulty based on the number of transfers in the reference route: routes with no transfers are labeled as easy, those with one transfer as medium, and all others as hard. To ensure balance, we set a fixed difficulty distribution threshold of $20:15:5$ (easy:medium:hard) for each map, generating $40$ questions. Once the quota for a difficulty level is reached on a given map, no additional questions of that level are retained.
@@ 3.1.3 Quality Control
To ensure the reliability and balance of the dataset, we perform quality control from three perspectives: correctness, diversity, and difficulty balance. Incorrect question–answer pairs are either manually corrected or discarded. We then involve both automatic checks and manual adjustments to ensure consistency and coverage across all difficulty levels. One reserved example is shown in Figure 1.
@@ 3.2 Dataset Statistics
The ReasonMap consists of $30$ high-resolution transit map images (see map sources in Appendix A.2) with an average resolution of $5,839\times 5,449$ pixels. In total, it contains $1,008$ question–answer pairs, including stop names in four languages (e.g., English, Hungarian, Chinese, and Italian). The distribution of question difficulty is as follows: $57.7\%$ are labeled as easy, $34.4\%$ as medium, and $7.8\%$ as hard. Additionally, a subset of $312$ samples is manually selected as the test set for the benchmark experiments described in Section 5, while the remaining samples serve as a training set for future use. To ensure diversity and difficulty balance, the test set includes $11$ cities with a $4:3:4$ map difficulty ratio and a question difficulty distribution ($181$ easy, $108$ medium, $23$ hard) that maintains consistency with the full dataset.
4 Evaluation Framework
This section systematically introduces a two-level evaluation framework for assessing model performance on the ReasonMap. This framework separately evaluates the correctness and quality of answers produced by models. Specifically, we quantify correctness using accuracy and design map score to measure the quality of answers, considering multiple factors (e.g., route efficiency and alignment with the reference routes from map services).
Preparation for Evaluation. We first parse the model-generated answers according to the required format. Answers that do not comply with the specified format or cannot be parsed due to model hallucination [58] are marked as invalid. Invalid responses are excluded from subsequent evaluations, with accuracy and map score set to zero. For the correctness evaluation, we utilize the Map Metadata mentioned in Section 3.1.1 as ground truth. For the quality evaluation, we adopt the collected reference routes as presented in Section 3.1.2 as the ground truth.

### v3

# ReasonMap: Towards Fine-Grained Visual Reasoning from Transit Maps

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated significant progress in semantic scene understanding and text-image alignment, with reasoning variants enhancing performance on more complex tasks involving mathematics and logic. However, their proficiency in tasks requiring both fine-grained visual understanding and spatial reasoning remains underexplored. To bridge this gap, we introduce ReasonMap, a novel benchmark specifically designed to evaluate these capabilities. ReasonMap encompasses high-resolution transit maps from $30$ cities and includes 1,0081{,}008 question-answer pairs spanning two question types and three templates. Furthermore, we design a two-level evaluation pipeline that properly assesses answer correctness and quality. Our comprehensive evaluation of $16$ popular MLLMs reveals a counterintuitive pattern: among open-source models, base variants outperform their reasoning-tuned counterparts, whereas the opposite trend is observed in closed-source models. Further analysis under the visual-masking setting confirms that strong performance necessitates direct visual grounding, rather than relying solely on language priors. We further establish a training baseline with reinforcement fine-tuning, providing a reference for future exploration. We hope this benchmark study offers new insights into visual reasoning and helps investigate the gap between open- and closed-source models.
Figure 1: Overview of ReasonMap. We present a novel benchmark tailored for evaluating the fine-grained visual reasoning capabilities of MLLMs. The dataset comprises 1,0081{,}008 question-answer pairs derived from high-resolution transit maps across $30$ cities in $13$ countries, featuring diversely structured questions. Further details on the dataset construction pipeline are provided in Section 3.
@@ 1 Introduction
Multimodal large language models (MLLMs) [1, 3, 102, 29, 42] have recently achieved notable advancements across a range of vision-language tasks, including visual grounding [50, 83, 80, 76], reasoning segmentation [9, 96, 53, 37, 70], and text-image alignment [88, 85]. Building upon these developments, reasoning MLLMs [47, 22, 64, 74, 49, 6, 51] have further improved performance on complicated visual reasoning tasks such as visual math problems [84, 67], visual question answering (VQA) [59, 87, 67], and spatial reasoning [59, 40, 11, 16]. These capabilities are critical for a wide range of real-world applications, including embodied AI and autonomous driving [12, 68, 10, 13]. As visual tasks grow in complexity and practical relevance, the need for rigorous benchmarks to assess complex visual reasoning becomes increasingly essential.
To address the growing demand for the evaluation of visual reasoning, several benchmarks have been proposed. Existing multimodal reasoning datasets, such as MathVQA [67], MMMU [87], and MathVerse [92], primarily assess symbolic or mathematical reasoning, where visual understanding plays a limited role. Conversely, datasets emphasizing fine-grained visual comprehension and retrieval like VisuLogic [79], VisualPuzzles [60], and VBench [75] require detailed perception but minimal planning or spatial reasoning. CityBench [13], DriveBench [76], and MapBench [77] take a step toward spatial reasoning while remain coarse in granularity. In contrast, our ReasonMap bridges these gaps by jointly evaluating fine-grained visual understanding, spatial reasoning and planning over high-resolution and information-dense transit maps.
@@
In this paper, we introduce ReasonMap (Figure 1), a benchmark designed to evaluate the fine-grained visual understanding and spatial reasoning capabilities of MLLMs using high-resolution transit maps. As structured and information-dense visual artifacts, maps inherently require precise spatial interpretation, making them well-suited for assessing detailed visual reasoning. ReasonMap comprises $1,008$ human-verified question-answer pairs spanning $30$ cities across $13$ countries. Each instance includes a map, two stops, two questions (short and long), multiple reference routes, and two difficulty labels (map and question difficulty). The questions cover two types and three question templates capturing both coarse and fine-grained spatial reasoning. To ensure data quality, we perform manual route verification, promote question diversity, and balance difficulty distribution. For evaluation, we propose a two-level framework that independently measures answer correctness (via accuracy) and quality (via a proposed map score), reflecting both feasibility and efficiency.
@@
We conduct comprehensive experiments on $16$ widely-used MLLMs, encompassing both base and reasoning models. Our results reveal a counterintuitive finding: among open-source models, base variants outperform their reasoning counterparts, whereas the opposite holds for closed-source models. Moreover, when only textual inputs are provided, models can still answer some questions based on inner knowledge, but in most cases, their performance noticeably drops. This highlights a critical limitation in the current model behavior. While some models can leverage prior knowledge and textual cues to solve certain tasks, fine-grained visual reasoning tasks requiring genuine understanding still necessitate effective integration of multimodal information for robust reasoning.
We further establish a training baseline with Group Relative Policy Optimization (GRPO) [57]. Guided by our evaluation framework, we design two reward components (i.e., accuracy reward and format reward) as reinforcement signals to optimize model behavior during training. Under a cross-city setting, where training and test maps are completely disjoint, the reinforcement fine-tuning consistently improves performance across models of different scales.
Our main contributions are summarized as follows: (1) We develop an extensible, semi-automated pipeline for dataset construction, facilitating scalable expansion to additional maps and cities ; (2) We propose a structured two-level evaluation framework that separately quantifies answer correctness and quality using accuracy and the proposed map score, respectively, enabling fine-grained answer assessment; (3) A comprehensive benchmarking study is conducted across $16$ MLLMs, providing insights into model performance, robustness, and the interplay between visual and textual cues, thereby informing future research on visual reasoning; and (4) We establish a training baseline with reinforcement fine-tuning as a reference for future research.
@@ 2 Related Work
Table 1: Comparison between ReasonMap and existing multimodal reasoning datasets. For entries in the dataset size column with notation like ($\times n$), each base problem has multiple versions to enforce visual grounding. Specifically, VGRP-Bench is constructed by sampling over 20 core puzzles; MathVerse generates six multimodal variants per problem with different levels of visual and textual information.
Name Year Dataset Size Avg. Resolution Training Set Step Evaluation Multilingual
MMMU [87] 2024 11.511.5k $684\times 246$
MathVerse [92] 2024 2,6122{,}612 ($\times 6$) $577\times 487$
VisuLogic [79] 2025 1,0031{,}003 $601\times 331$
VisualPuzzles [60] 2025 1,1681{,}168 $767\times 464$
VGRP-Bench [52] 2025 2020 ($\times 5$) $790\times 790$
R-Bench [23] 2025 $665$ $629\times 348$
@@
ReasonMap 2025 1,0081{,}008 ($\times 2$) $5{,}839\times 5{,}449$
Reasoning in LLMs & MLLMs. Recent advances in large language models (LLMs) have demonstrated significant improvements in reasoning capabilities through reinforcement fine-tuning paradigms [47, 22, 15, 26], which leverage GRPO [57] to unlock the reasoning potential of LLMs. This paradigm has also been extended to the multimodal domain, with increasing interest in applying reinforcement learning (RL) to visual reasoning [65, 35, 44, 61, 58]. Both open-source and closed-source communities have introduced advanced reasoning MLLMs built upon earlier systems [102, 3, 82, 48]. Notable open-source models include Kimi-VL [64], Skywork-R1V [74, 49], and Qwen-QvQ [51], whereas Doubao-1.5-Pro [6], Seed1.5-VL [21], OpenAI o3 [48], OpenAI 4o [46], and Gemini [19] represent leading closed-source efforts. Despite recent progress, systematic evaluation of fine-grained visual reasoning in MLLMs remains limited, as existing benchmarks primarily target coarse-grained tasks and fail to capture model performance on complex real-world visual content.
Multimodal Reasoning Datasets. As multimodal reasoning has rapidly progressed, various benchmarks have emerged to evaluate MLLMs across different reasoning dimensions (see summary in Table 1). Datasets such as VBench [75], VisualPuzzles [60], VisuLogic [79], and VGRP-Bench [52] primarily examine abstract visual reasoning through synthetic tasks involving logic, structure, and pattern recognition. In parallel, MapBench [77], CityBench [13], and DriveBench [76] shift focus to real-world spatial reasoning, assessing model performance on urban and autonomous driving scenarios. For mathematical reasoning, benchmarks like MathVQA [67], MMMU [87], and MathVerse [92] integrate multimodal inputs, with MathVerse notably introducing varied problem formats to strengthen visual dependence. Geospatial benchmarks [36, 81, 86] have also been proposed to evaluate spatial reasoning and map understanding capabilities. Unlike previous benchmarks that primarily emphasize either fine-grained visual understanding (e.g., VBench) or spatial reasoning (e.g., MapBench), our ReasonMap benchmark jointly evaluates both aspects with high-resolution transit maps.
Map-based Spatial Reasoning. Among the many directions of multimodal reasoning, map-based spatial reasoning has emerged as a crucial area, with broad applications in navigation, urban planning, and autonomous systems [5, 54, 78, 69, 71]. Recent efforts have focused on enabling models to interpret and reason over various types of map data. CityBench [13] provides a dataset for evaluating urban scene understanding, while MapLM [7] introduces a benchmark for map and traffic scene understanding. PlanAgent [100] and PReP [90] explore embodied planning in environments that require interpreting map information. MapEval [11] proposes a structured evaluation suite for map reasoning, and GeoNav [78] investigates geospatial navigation using LLMs. Most existing methods [11, 13, 100] depend on external tools (e.g., map services or APIs) to complete spatial tasks, which often bypasses the need for genuine visual reasoning. However, spatial reasoning based on visual understanding remains essential. Our work aims to fill this gap by evaluating such capabilities without tool assistance.
@@ 3 ReasonMap Construction
In this section, we first present the complete dataset building pipeline as shown in Figure 2, which consists of three main stages: (1) data collection and preprocessing, (2) question–answer pair construction, and (3) quality control. We then report comprehensive statistics of the dataset.
@@ 3.1.1 Data Collection and Preprocessing
We collect and manually select $30$ high-resolution transit maps covering $30$ cities across $13$ countries from publicly available online sources, in compliance with relevant licenses and regulations, ensuring diversity and a balanced range of map difficulty. We then leveraged MLLMs (e.g., GPT-4o) to extract the names of transit lines and their corresponding stops, followed by manual correction to ensure correctness. Special cases like transfer stops and branch-starting stops were annotated in a standardized format appended to the respective stop entries. Finally, for subsequent usage, all route and stop information was saved in a unified JSON format, referred to as the Metro Data.
@@ 3.1.2 Question-Answer Pair Construction
The construction of question–answer pairs involves three key steps: (1) Question Generation, where we formulate questions based on predefined templates; (2) Reference Route Collection, where we obtain corresponding reference routes using Google Map and Gaode Map; and (3) Label Annotation, where we properly assign difficulty labels for both maps and questions.
@@
Label Annotation. Two levels of difficulty labeling are included in this stage. For map difficulty, we assign each map to one of three levels (easy, medium, hard) based on the number of transit lines and transfer stations, ensuring a balanced split of $10$ maps per level. For question difficulty, we assign difficulty based on the number of transfers in the reference route: routes with no transfers are labeled as easy, those with one transfer as medium, and all others as hard. To ensure balance, we set a fixed difficulty distribution threshold of $20:15:5$ (easy:medium:hard) for each map, generating $40$ questions. Once the quota for a difficulty level is reached on a given map, no additional questions of that level are retained. Additionally, we provide a more fine-grained taxonomy of questions in Appendix A.2.
@@ 3.1.3 Quality Control
To ensure the reliability and balance of the dataset, we perform quality control from three perspectives: correctness, diversity, and difficulty balance (more details in Appendix A.3). Incorrect question–answer pairs are either manually corrected or discarded. We then involve both automatic checks and manual adjustments to ensure consistency and coverage across all difficulty levels. One reserved example is shown in Figure 1.
@@ 3.2 Dataset Statistics
The ReasonMap consists of $30$ high-resolution transit map images (see map sources in Appendix A.4) with an average resolution of $5,839\times 5,449$ pixels. In total, it contains $1,008$ question–answer pairs. The distribution of question difficulty is as follows: $57.7\%$ are labeled as easy, $34.4\%$ as medium, and $7.8\%$ as hard. Additionally, a subset of $312$ samples is manually selected as the test set for the benchmark experiments described in Section 5, while the remaining samples serve as a training set. To ensure diversity and difficulty balance, the test set includes $11$ cities with a $4:3:4$ map difficulty ratio and a question difficulty distribution ($181$ easy, $108$ medium, $23$ hard) that maintains consistency with the full dataset. Moreover, ReasonMap includes inter-modal transfers in cities like Sydney, where subways, light rail, and airport lines converge.
4 Evaluation Framework & Training Baseline
In this section, we first present a two-level evaluation framework for systematically assessing model performance, which separately measures the correctness and quality of model-generated answers. We further introduce a GRPO-based training baseline with carefully designed rewards.
Preparation for Evaluation. We first parse the model-generated answers according to the required format. Answers that do not comply with the specified format or cannot be parsed due to model hallucination [4] are marked as wrong answer. For the correctness evaluation, we utilize the Metro Data mentioned in Section 3.1.1 as ground truth. For the quality evaluation, we adopt the collected reference routes as presented in Section 3.1.2 as the ground truth.